What is the formula weight for iron(III) oxide, in amu?
Answer #1

iron(III) oxide

Fe2O3

Atomic mass of Fe = 55.84 amu

atomic mass of oxygen = 16 amu

formula weight of Fe2O3 = 2*55.84 +3*16

                                       = 111.68 + 48 = 159.68 amu